Two blocks of mass and M connected by a massless spring of spring con stant & are placed on a smooth horizontal table. Determine the equations of




motion using Newtonian mechanics

1
Expert's answer
2022-01-02T18:23:15-0500

Answer

The kinetic energy and potential energy from the coupling spring constant k for this system can written as

"T= \\frac1{2} m \nx\n\u02d9\n \n2\n + \\frac1{2} M \ny\n\u02d9\n\u200b\t\n \n2"

And

"V=\frac1{2}k(x−y) ^2"


So lagrangial of this system can witten as

"L=\\frac1{2} m \nx\n\u02d9\n \n2\n + \\frac1{2} M \ny\n\u02d9\n\u200b\t\n \n2-\\frac1{2}k(x\u2212y) ^2"

Now calculating as below

"\\frac{\\partial L}{\\partial \\dot{x}}=m\\dot{x}, \\\\ \\frac{\\partial L }{\\partial x}=-k(x-y)\\\\ \n \\frac{\u2202L}{\u2202 \nx}\n\u02d9\n =m \nx\n\u02d9\n ,\\\\\u2202L\/\u2202x=\u2212k(x\u2212y)\\\\ \n\n\\partial L \/ \\partial \\dot{y}=M \\dot{y} , \\\\ \\partial L\/ \\partial y=k(x-y)\\\\\u2202L\/\u2202 \ny\n\u02d9\n\u200b\t\n =M"

Now lagrangian equation of motion is given by

"\frac{d}{dt}{\frac{\partial L}{ \partial \dot{x}}}-\frac{\partial L}{\partial x}=0, \\"

"\frac{d}{dt}{\frac{\partial L}{ \partial \dot{y}}}-\frac{\partial L}{\partial y}=0,"

Putting all values then we get

Equation of motion

"m\ddot{x}+k(x-y)=0\\m\ddot{y}+k(y-x)=0"